# Logic Jailbreak: Efficiently Unlocking LLM Safety Restrictions Through Formal Logical Expression

### v1

@@ Abstract
Despite substantial advancements in aligning LLMs with human values, current safety mechanisms remain susceptible to jailbreak attacks. We hypothesize that this vulnerability stems from distributional discrepancies between alignment-oriented prompts and malicious prompts. To investigate this, we introduce LogiBreak—a novel and universal black-box jailbreak method that leverages logical expression translation to circumvent LLM safety systems. By converting harmful natural language prompts into formal logical expressions, LogiBreak exploits the distributional gap between alignment data and logic-based inputs, preserving the underlying semantic intent and readability while evading safety constraints. We evaluate LogiBreak on a multilingual jailbreak dataset spanning three languages, demonstrating its effectiveness across various evaluation settings and linguistic contexts. The code is available at https://anonymous.4open.science/r/Logibreak-DEBF. Warning: This paper contains potentially harmful text.
@@
Jingyu Peng‡§  Maolin Wang§, Nan Wang, Xiangyu Zhao§ 11footnotemark: 1, Jiatong Li, Kai Zhang, Qi Liu 11footnotemark: 1 University of Science and Technology of China, § City University of Hong Kong, Universiteit van Amsterdam jypeng28@mail.ustc.edu.cn morin.wang@my.cityu.edu.hk
Figure 1: A Venn diagram illustrating the hypothesis that model vulnerabilities arise from distributional differences between alignment and malicious prompts, highlighting how special token prompts, low-resource languages (LRL) and multilingual prompts are more prone to jailbreak.
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive capabilities and are widely used across industry and research, with examples including ChatGPT OpenAI (2023), DeepSeek Liu et al. (2024), and Llama Grattafiori et al. (2024). To mitigate misuse, LLMs are typically aligned with human values through post-training methods Zhou et al. (2024b). However, these safeguards can be bypassed by carefully crafted inputs—so-called“jailbreak” attacks Yi et al. (2024)—posing serious risks to safe and responsible AI deployment.
@@
Jailbreak attacks are broadly categorized into white-box and black-box methods Das et al. (2025). White-box approaches use model internals like gradients or logits to find adversarial prompts Chao et al. (2023); Li et al. (2025), but are impractical for closed-source models. Black-box methods, which only require API access, use optimization algorithms Jawad and BRUNEL (2024); Basani and Zhang (2024) or crafted instructions Shen et al. (2024); Liu et al. (2023); Zhou et al. (2024a); Yu et al. (2024). These techniques are often inefficient or brittle, as they rely on specific prompt patterns and are easily disrupted by model updates or system prompt changes. This motivates a more fundamental understanding of alignment vulnerabilities.
@@
We propose a simple yet general hypothesis: the vulnerability of aligned language models arises from token-level distributional differences between alignment and malicious prompts (Figure 1). Alignment training typically covers a narrow range of token sequences—discrete units like subwords or characters—without regard to meaning. As a result, prompts with unfamiliar token patterns may fall outside the model’s aligned behavior, even if their semantic intent remains harmful.
This gap between token-level form and semantic intent allows adversarial prompts to bypass safeguards. Malicious inputs can be rephrased or encoded to preserve meaning while presenting token distributions that appear benign or novel to the model. Empirical findings support this: attacks are more successful in low-resource languages Yong et al. (2023), with multilingual prompts Deng et al. , and through the use of special tokens Yu et al. (2024)—all strategies that shift token distributions without altering intent.
Under this hypothesis, we aim to find a universal transformation that shifts natural language prompts into an unaligned token space while preserving semantics interpretable by LLMs. Translating prompts into logical expressions—particularly first-order logic (NL-FOL)—offers a principled approach, maintaining semantic integrity in a structurally distinct form. While NL-FOL plays a foundational role in many logic-driven NLP tasks Bos and Markert (2005); Lee et al. (2025); Angeli and Manning (2014); Xia et al. (2025); Guu et al. (2015); Hamilton et al. (2018), its alignment-related implications remain underexplored. Prior work on cipher-based methods (e.g., Base_64, Caesar) Yuan et al. ; Yong et al. (2023) also shifts token distributions but often suffers from generation failures, and poor readability, limiting practical use, as shown in 3.5.
In this paper, we propose a novel jailbreak method callded LogiBreak, that exploits logical expressions to circumvent LLM safety guardrails. Our approach systematically translates harmful or prohibited natural language prompts into formal logical representations that preserve the semantic content while evading detection by alignment mechanisms as shown in Figure 2. By leveraging the distributional gap between alignment training data and logical formulations, this jailbreak technique enables consistent bypassing of safety filters without requiring white-box access to model parameters. Importantly, our method is language-agnostic and universally applicable across multiple languages, providing a more generalizable jailbreak method than previous approaches.
Moreover, LogiBreak’s effectiveness reveals a key limitation of current safety alignment methods, which largely focus on token-level defenses while neglecting semantic vulnerabilities. Although LogiBreak operates at the token level, its transformations preserve semantic alignment with the original prompts, exposing a critical gap in existing safeguards. This underscores the need for more comprehensive safety measures that address deeper semantic understanding rather than relying solely on surface-level defenses.
@@
Figure 2: t-SNE visualization of the raw requests and their corresponding logically reformed versions.
@@
  • We present a general and principled hypothesis that the vulnerability of aligned LLMs arises from distributional disparities between alignment prompts and adversarial prompts—a perspective substantiated by many existing jailbreak methods.
  • We introduce LogiBreak, a general and effective black-box jailbreak method that systematically translates malicious prompts into logical expressions, preserving their semantic content while shifting the prompt distribution away from alignment-trained data.
  • We construct a novel multilingual logical expression jailbreak dataset based on JailbreakBench, comprising 100 malicious requests expressed as logical formulas in English, Chinese, and Dutch—filling the current gap in benchmarks for evaluating LLMs robustness against logical expression–based attacks.
  • LogiBreak’s ability to exploit distributional shifts without altering semantic exposes a core weakness in current LLM safety alignment: a reliance on surface-level cues while overlooking semantic consistency. This reveals a deeper structural vulnerability in existing defense mechanisms.
@@ 2.1 Task Definition
A jailbreak task involves crafting prompts that cause LLMs to respond to harmful requests that they would typically refuse to answer. These requests often belong to predefined categories recognized as harmful by model providers—such as those related to misinformation, violence, hate speech, and other prohibited content.
We begin by formally defining the jailbreak task in the context of a target LLM $\mathcal{M}$, which maps an input prompt $x\in\mathcal{X}$ to an output $y=\mathcal{M}(x)\in\mathcal{Y}$.
Let $\mathcal{X}_{\text{harmful}}\subset\mathcal{X}$ denote a predefined set of direct harmful requests, such as prompts involving violence, misinformation, or other categories typically restricted under safety policies.
Given that modern LLMs are commonly aligned using large-scale datasets that explicitly include such harmful examples for the purpose of safety fine-tuning, we posit the following inclusion relationship:
@@
where $\mathcal{D}_{\text{aligned}}$ represents the empirical data distribution used during the alignment stage. This assumption reflects the fact that direct harmful prompts are likely to have been encountered and explicitly refused during training.
Accordingly, a safety-aligned model $\mathcal{M}$ is expected to respond to any harmful input $x_{\text{harmful}}\in\mathcal{X}_{\text{harmful}}$ with a refusal output:
@@
where $\mathcal{Y}_{\text{refuse}}\subset\mathcal{Y}$ denotes the set of outputs indicating refusal to comply (e.g., "I’m sorry, but I can’t help with that.").
The objective of the jailbreak task is to construct a transformation function $\mathcal{F}$ such that, given a harmful prompt $x_{\text{harmful}}$, the transformed version $x^{\prime}_{\text{harmful}}=\mathcal{F}(x_{\text{harmful}})$ satisfies:
$\mathcal{M}(x^{\prime}_{\text{harmful}})\notin\mathcal{Y}_{\text{refuse}}.$
In other words, the transformed prompt successfully circumvents the model’s safety mechanisms and elicits a response that would not be produced for the original input.
Crucially, the transformation $\mathcal{F}$ must also preserve the semantic intent of the original harmful prompt. This constraint ensures that the jailbreak does not simply obfuscate the prompt, but instead maintains its intended meaning. Formally, we require:
@@
where $\text{Sim}(\cdot,\cdot)$ is a semantic similarity function and $\tau$ is a predefined threshold enforcing sufficient preservation of meaning.
@@ 2.2 Approach
As shown in Figure 3, to construct an effective transformation function $\mathcal{F}$ that both shifts prompts out of the alignment training distribution and preserves their semantic intent, we leverage natural language to logical expression translation. Formally, the transformation function can be defined as:
$\mathcal{F}:\mathcal{X}_{\text{harmful}}\rightarrow\mathcal{X}_{\text{logic}},$
Since the safety of logical expressions are never considered in previous works on safety alignment, we assume that the translated prompts lie outside the support of $\mathcal{D}_{\text{aligned}}$, i.e.,
@@
Notably, unlike natural language translation, this approach is language-agnostic and can be applied across diverse languages, consistently preserving semantic content regardless of the source language, as illustrated in the bottom-left portion of Figure 3. Moreover, from a linguistic perspective, this translation process naturally preserves the semantic intent of the original prompt, as logical forms are designed to capture meaning independent of surface linguistic variation.
Given that the task of translating natural language into logical forms has been shown to be tractable for LLMs Lee et al. (2025); Xiong et al. (2024); Lalwani et al. (2024), we can leverage LLMs themselves to automatically generate logical expressions from natural language inputs.
@@
[…]
@@
$x^{\prime}=x_{\text{context}}||x_{\text{logic}}||x_{\text{instruct}}$
The complete prompts utilized in LogiBreak are presented in Section B.
Table 1: LogiBreak jailbreak performance on six LLMs evaluated with three distinct methods.
Rule_Judge Llama3-70B_Judge GPT4-Judge
@@ 3.1 Dataset Construction
To evaluate the multilingual effectiveness of LogiBreak, we construct a novel logical expression jailbreak dataset based on JailbreakBench Chao et al. . This effort is motivated by a current gap in existing benchmarks: while many datasets evaluate LLM robustness against traditional jailbreaks, there is a lack of systematic resources specifically targeting logical expression–based attacks. Our dataset aims to fill this gap and provide a more rigorous testbed for assessing the resilience of LLMs to this underexplored jailbreak.
The dataset comprises 100 malicious requests, each manually translated into three distinct languages: English, Chinese, and Dutch. Examples of both the natural language prompts and their corresponding logical expressions are illustrated in Figure 3.
Specifically, we first utilize GPT-3.5 to translate the original English prompts into Chinese and Dutch. However, due to the model’s built-in safety mechanisms, it frequently refuses to process harmful content, resulting in incomplete or deflected translations. To address this issue, we supplement the translation process with conventional machine translation tools and subsequently engage native speakers to refine and validate the final outputs. This multi-stage process ensures both the semantic fidelity and linguistic accuracy of the malicious requests across all three languages, thereby enhancing the overall quality and reliability of the dataset.
Next, the logical expression corresponding to each request is generated using GPT-3.5, following the prompt template illustrated in Section A. Due to the safety constraints of the model, a small number of translation attempts were refused during this process. To improve coverage, we employed additional models—GPT-4o and DeepSeek V3—to reattempt translation in cases of failure. For prompts that remained unprocessed after these steps, we enlisted a PhD student specializing in linguistics to manually complete the logical expression translation.
Table 2: Performance comparison with five black-box jailbreak baselines measured by ASR@5.

### v2

@@ Abstract
Despite substantial advancements in aligning LLMs with human values, current safety mechanisms remain susceptible to jailbreak attacks. We attribute this vulnerability to the distributional discrepancies between alignment-oriented prompts and malicious prompts. To investigate this, and drawing inspiration from logic-driven NLP tasks, we introduce LogiBreak, a universal black-box jailbreak method that utilizes logical expression translation to bypass LLM safety mechanisms. By converting harmful natural language prompts into formal logical expressions, LogiBreak exploits the distributional gap between alignment data and logic-expressed inputs, preserving the underlying semantic intent and readability while evading safety constraints. Furthermore, to fill the gap of existing benchmarks that lack systematic resources specifically targeting logical expression-based attacks against LLM robustness, we construct a novel multilingual logical expression jailbreak dataset for evaluation. Our evaluations of LogiBreak in five languages demonstrate its effectiveness and generalizability in various linguistic contexts. The code is available at https://anonymous.4open.science/r/Logibreak-DEBF.
Warning: This paper contains potentially harmful text.
@@
Jingyu Peng‡§, Maolin Wang§, Nan Wang, Jiatong Li, Yuchen Li, Yuyang Ye§, Wanyu Wang§, Pengyue Jia§, Kai Zhang11footnotemark: 1, Xiangyu Zhao§11footnotemark: 1 University of Science and Technology of China, § City University of Hong Kong, Universiteit van Amsterdam, Baidu Inc. jypeng28@mail.ustc.edu.cn, morin.wang@my.cityu.edu.hk
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive capabilities and are widely used across industry and research, with examples including ChatGPT OpenAI (2023), DeepSeek Liu et al. (2024a), and Llama Grattafiori et al. (2024). To mitigate misuse, LLMs are typically aligned with human values through post-training methods Zhou et al. (2024b). However, these safeguards can be bypassed by carefully crafted inputs, so-called "jailbreak" attacks Yi et al. (2024), posing serious risks to the deployment of safe and responsible AI.
@@
Jailbreak attacks are broadly categorized into white-box and black-box methods Das et al. (2025). White-box approaches use model internals like gradients or logits to generate adversarial prompts Chao et al. (2023); Li et al. (2025), but are impractical for closed-source models. Black-box methods, which only require API access, use optimization algorithms Jawad and BRUNEL (2024); Basani and Zhang (2024) or crafted instructions Shen et al. (2024); Liu et al. (2023); Zhou et al. (2024a); Yu et al. (2024). These techniques are often inefficient or brittle, as they rely on specific prompt patterns and are easily disrupted by model updates or system prompt changes Xie et al. (2023); Liu et al. (2024c). This motivates a more fundamental understanding of the alignment vulnerabilities of jailbreak attacks.
Figure 1: A Venn diagram illustrating the hypothesis that model vulnerabilities arise from distributional differences between alignment and malicious prompts, highlighting how special token prompts, low-resource languages (LRLs) and multilingual prompts are more prone to jailbreak.
@@
In pursuing this goal, as shown in Figure 1, we propose a simple yet general hypothesis: the vulnerability of aligned language models arises from token-level distributional differences between alignment and jailbreak prompts. Alignment training typically covers a narrow range of token sequences, which are discrete units like words, without regard to semantic meaning. As a result, prompts with unfamiliar token patterns may fall outside the model’s aligned behavior, even if their semantic intent remains harmful. This gap between token-level form and semantic intent allows adversarial prompts to bypass safeguards. Malicious inputs can be transformed to preserve meaning while presenting token distributions that appear benign or novel to the model.
Empirical evidence supports this hypothesis: attacks tend to be more effective on low-resource languages (LRLs) Yong et al. (2023), when using multilingual prompts Deng et al. (2023), special tokens Yu et al. (2024), or cipher-based techniques (e.g., Base_64, Caesar) Yuan et al. (2024); Yong et al. (2023). These methods commonly aim to shift token distributions without changing the underlying intent.
However, these approaches face notable limitations. Language-based methods are constrained by the limited number of available languages and cannot ensure that LRLs are entirely excluded from alignment data. Techniques involving special tokens or ciphers often struggle to preserve the original semantic meaning. Moreover, the ineffectiveness of crafting instructions can also be attributed to their failure to sufficiently transform the malicious request, since they only add a prefix or suffix without modifying the raw request.
Along this hypothesis, we aim to find a universal transformation that shifts natural language prompts into an unaligned token space while preserving semantics interpretable by LLMs. Translating prompts into logical expressions, particularly first-order logic (NL-FOL), offers a principled approach. Under the theoretical framework of formal semantics, an ideal translation from NL to FOL constitutes a truth-preserving mapping that maintains full compositional meaning, where the logical form is isomorphic to the semantic content of the natural language utterance Winter (2016). Practically, the back-translation experiments in Table 1 validate the semantic consistency of the NL-FOL translation, achieving a satisfactory BERTScore F1 of approximately 0.9 across all five languages.
English Chinese Dutch Japanese Spanish
Precision 0.948 0.921 0.885 0.914 0.924
Recall 0.944 0.941 0.912 0.925 0.903
F1 0.946 0.931 0.898 0.919 0.913
Table 1: Evaluation of semantic consistency between natural language and FOL by back-translating FOL to natural language and evaluating with BERTScore.
From another perspective, while NL-FOL plays a foundational role in many logic-driven NLP tasks Bos and Markert (2005); Lee et al. (2025); Angeli and Manning (2014); Xia et al. (2025); Guu et al. (2015); Hamilton et al. (2018), its safety-related implications remain underexplored.
@@
Figure 2: t-SNE visualization of raw requests and their corresponding reformed versions in English. Blue points represent raw requests, while red points denote their reformed counterparts.
In this paper, we propose a black-box jailbreak method called LogiBreak, which exposes a core weakness in current LLM safety alignment: a reliance on token-level patterns rather than true semantic understanding. Our method systematically translates harmful or prohibited prompts into logical expressions that preserve their meaning while evading alignment-based detection, as shown in Figure 2. By leveraging the distributional gap between alignment training data and logical formulations, LogiBreak consistently bypasses safety filters across multiple languages.
Moreover, LogiBreak further reveals a key limitation of current safety alignment methods, which largely focus on token-level defenses while neglecting semantic vulnerabilities. Although LogiBreak operates at the token level, its transformations preserve semantic alignment with the original prompts, exposing a critical gap in existing safeguards. This underscores the need for more comprehensive safety measures that address deeper semantic understanding rather than relying solely on token-level defenses.
@@
  • We propose a principled hypothesis that LLM jailbreak vulnerability stems from distributional disparities between alignment and adversarial prompts, supported by evidence from existing attacks.
  • We present LogiBreak, a black-box jailbreak method that converts malicious prompts into logical expressions, preserving semantics while shifting prompt distribution.
  • We construct a multilingual logical-jailbreak dataset (English, Chinese, Dutch, Japanese, Spanish) based on JailbreakBench, filling a gap in evaluating LLM robustness to logic-based attacks.
  • LogiBreak’s success exposes a core weakness in current LLM safety alignment: a reliance on token-level cues while overlooking semantic consistency. This reveals a deeper structural vulnerability in existing defenses.
@@ 2.1 Task Definition
A jailbreak task involves crafting prompts that cause LLMs to respond to harmful requests that they would typically refuse to answer. These requests often belong to predefined categories recognized as harmful by model providers.
Formally, consider an LLM $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$, mapping an input prompt $x\in\mathcal{X}$ to an output $y=\mathcal{M}(x)\in\mathcal{Y}$. Let $\mathcal{X}_{\text{harmful}}\subset\mathcal{X}$ denote a set of harmful prompts.
Given that modern LLMs are commonly aligned using large-scale datasets that contain such harmful examples for the purpose of safety alignment, we posit the following inclusion relationship:
@@
reflecting that the model has been exposed to such prompts during training and is fine-tuned to refuse them. Hence, a safety-aligned model is expected to respond to any $x_{\text{harmful}}\in\mathcal{X}_{\text{harmful}}$ with outputs indicating refusal:
@@
where $\mathcal{Y}_{\text{refuse}}\subset\mathcal{Y}$ contains refusal responses like “I’m sorry, but I can’t assist with that.”
The objective of the jailbreak task is to construct a transformation $\mathcal{F}$ that satisfy:
$x^{\prime}_{\text{harmful}}=\mathcal{F}(x_{\text{harmful}}),\quad\mathcal{M}(x^{\prime}_{\text{harmful}})\notin\mathcal{Y}_{\text{refuse}}.$
In other words, the transformed prompt $x^{\prime}_{\text{harmful}}$ successfully circumvents the model’s safety mechanisms and elicits a response that would not be produced for the original input.
Crucially, the transformation $\mathcal{F}$ must preserve the semantic intent of the original prompt. This constraint ensures that the jailbreak does not simply obfuscate the prompt, but instead maintains its intended meaning. Formally, we require:
@@
where $\text{Sim}(\cdot,\cdot)$ is a semantic similarity measure and $\tau$ is a threshold enforcing preservation of the original intent.
In summary, the jailbreak task can be formulated as the following constrained optimization:
$\max_{\mathcal{F}}$ $\Pr\left[\mathcal{M}(x^{\prime}_{\text{harmful}})\notin\mathcal{Y}_{\text{refuse}}\right]$
s.t. $x^{\prime}_{\text{harmful}}=\mathcal{F}(x_{\text{harmful}}),$
$\text{Sim}(x_{\text{harmful}},x^{\prime}_{\text{harmful}})\geq\tau.$
@@ 2.2 Approach
Existing jailbreak methods, however, struggle with this constrained optimization. They either fail to effectively avoid the aligned prompt distribution or lack consistent semantic preservation, which ultimately limits their effectiveness.
Inspired by the technique of NL-FOL translation that preserves meaning but has unexamined safety vulnerabilities in the context of LLMs, Our approach uses NL-FOL translation to construct $\mathcal{F}$ as illustrated in Figure 3. Formally, the transformation function can be defined as:
$\mathcal{F}:\mathcal{X}_{\text{harmful}}\rightarrow\mathcal{X}_{\text{logic}}.$
Since the safety of logical expressions has not been considered in prior work on safety alignment, we reasonably assume that the translated prompts lie outside the support of $\mathcal{D}_{\text{aligned}}$, i.e.,
@@
This is a reasonable working assumption based on our survey of prior alignment datasets, which focus on natural-language token distributions rather than formal logic representations.
Notably, unlike natural language translation, this approach is language-agnostic and can be applied across diverse languages, consistently preserving semantic content regardless of the source language, as shown in the bottom-left portion of Figure 3. Linguistically, translating prompts into logical forms preserves their semantic intent by capturing meaning independent of surface language variations. This claim is empirically validated by the back-translation results in Table 1, which show high semantic consistency.
Given that the task of translating natural language into logical forms has been shown to be tractable for LLMs Lee et al. (2025); Xiong et al. (2024); Lalwani et al. (2024), we can even leverage LLMs to automatically generate logical expressions from natural language inputs.
@@
$x_{\text{context}}=\text{``I am working on formal semantics.''}.$
@@
[…]
The complete prompts utilized in LogiBreak are presented in the Appendix.
Rule_Judge LLaMA_Judge GPT_Judge
@@
Table 2: LogiBreak jailbreak performance on six LLMs evaluated with three judges (in %).
@@ 3.1 Dataset Construction
To evaluate the effectiveness of LogiBreak across multiple languages, we construct a new logical expression jailbreak dataset based on JailbreakBench Chao et al. (2024), covering English, Chinese, Dutch, Japanese, and Spanish. Existing benchmarks primarily evaluate LLM robustness against traditional jailbreaks and lack systematic resources for logical-expression-based attacks. Our dataset addresses this gap and provides a more rigorous testbed for assessing model resilience. The five languages were chosen for both their broad linguistic coverage (reflecting major speaker demographics) and the research team’s proficiency, ensuring data integrity and reliable evaluation.
The dataset contains 100 malicious requests, first translated into the five languages using machine translation tools, as modern LLMs tend to refuse harmful content or produce unreliable translations. Each request was then converted into a logical expression using GPT-3.5 OpenAI (2023) in a few-shot setting, following the template provided in the Appendix B.
To ensure consistent quality, experienced bilingual volunteers refined and validated all translations. For each language, three PhD candidates in Logic or Linguistics (native speakers or CEFR C2-level) independently assessed semantic fidelity. Inter-annotator agreement exceeded Fleiss’ $\kappa\geq 0.90$ across languages, and remaining discrepancies were resolved through consensus discussions supervised by a senior logician. This two-stage process guarantees both semantic equivalence and linguistic authenticity, significantly enhancing the dataset’s reliability.
3.2 Evaluation
To evaluate the effectiveness of our jailbreak method, we measure the Attack Success Rate (ASR), where ASR@N denotes the success rate over N independent attack trials following previous works Andriushchenko and Flammarion (2025). To provide a comprehensive assessment of Logibreak’s performance, we adopt three complementary evaluation methods, following the approach of Andriushchenko and Flammarion (2025).
The first judge is a rule-based evaluation proposed by Zou et al. (2023), which identifies specific words or phrases that indicate the target language model has refused to respond. While useful, this approach is limited because it ignores the semantic content of the response, which is essential for determining whether the output is potentially harmful. To address this shortcoming, we additionally adopt the semantic evaluation protocols from Chao et al. (2023) and Chao et al. (2024), using GPT-4 and LLaMA-3 70B to assess outputs based on the same prompts employed in those studies.
For the additional four languages, we translated the corresponding prompts using GPT-3.5, with all translations manually reviewed to ensure semantic accuracy and contextual appropriateness. The implementation of all evaluations are included in the Appendix B.
3.3 Baselines
To evaluate the effectiveness of our proposed approach, we compare it against six SOTA black-box jailbreak methods from the literature: Caesar Yuan et al. (2024), PastTense Andriushchenko and Flammarion (2025), LRL-Comb Yong et al. (2023), AIM, Prefix Wei et al. (2023), and RefComp Wei et al. (2023). Detailed descriptions of these methods are provided in Appendix C.
3.4 Models
We evaluate our method across a diverse set of target models, encompassing both open-source and closed-source LLMs. The open-source models include Qwen-2.5-7B Yang et al. (2024), LLaMA-3-8B Grattafiori et al. (2024), DeepSeek V3 Liu et al. (2024a), and DeepSeek R1 Guo et al. (2025), while the closed-source models consist of GPT-3.5 OpenAI (2023) and GPT-4o-mini OpenAI (2024).
For all models, we adopt a black-box setting, wherein we interact with the models exclusively through their APIs, which ensures a fair comparison and allows us to evaluate the generalizability of LogicBreak across diverse architectures under the most stringent conditions.
3.5 Overall Results
The empirical results in Table 2 show that LogiBreak consistently improves ASR across various evaluation frameworks and judging models. Specifically, it achieves average ASR@5 gains of over 4.9×, 35.3×, and 17.8× under rule-based, LLaMA, and GPT judges, respectively, demonstrating the robustness and generalizability of logically guided transformations in bypassing safety constraints. In particular, even in the stricter ASR @ 1 setting, where only the first adversarial attempt counts, LogiBreak maintains an average ASR exceeding 40% across all judges, suggesting that a single transformation often suffices. Compared to baselines, LogiBreak outperforms all other methods across all LLMs and evaluations, as shown in Table 3. It is noteworthy that although Caesar achieved a 100% ASR under the rule-based judge, its ASR was significantly lower under other semantic judges. This indicates that the LLM failed to perform the encoding and decoding tasks accurately, resulting in outputs that lack meaningful content.
To confirm that LogiBreak consistently shifts malicious prompts into a distinct embedding space, we provide an extended t-SNE visualization in Appendix A.1.